# ResAD: Normalized Residual Trajectory Modeling for End-to-End Autonomous Driving

### v1

@@ Abstract
End-to-end autonomous driving (E2EAD) systems, which learn to predict future trajectories directly from sensor data, are fundamentally challenged by the inherent spatio-temporal imbalance of trajectory data. This imbalance creates a significant optimization burden, causing models to learn spurious correlations instead of causal inference, while also prioritizing uncertain, distant predictions, thereby compromising immediate safety. To address these issues, we propose ResAD, a novel Normalized Residual Trajectory Modeling framework. Instead of predicting the future trajectory directly, our approach reframes the learning task to predict the residual deviation from a deterministic inertial reference. The inertial reference serves as a counterfactual, forcing the model to move beyond simple pattern recognition and instead identify the underlying causal factors (e.g., traffic rules, obstacles) that necessitate deviations from a default, inertially-guided path. To deal with the optimization imbalance caused by uncertain, long-term horizons, ResAD further incorporates Point-wise Normalization of the predicted residual. It re-weights the optimization objective, preventing large-magnitude errors associated with distant, uncertain waypoints from dominating the learning signal. Extensive experiments validate the effectiveness of our framework. On the NAVSIM benchmark, ResAD achieves a state-of-the-art PDMS of 88.6 using a vanilla diffusion policy with only two denoising steps, demonstrating that our approach significantly simplifies the learning task and improves model performance. The code will be released to facilitate further research.
@@
Figure 1: (a) Comparison of trajectory distributions under different modeling strategies. Raw trajectories exhibit significant mean drift and increasing variance over the prediction horizon. Trajectory Residual Modeling centers the distribution around zero. Point-wise Residual Normalization further stabilizes variance for a simpler learning objective. (b) Comparisons between existing methods and the proposed ResAD. Instead of predicting the trajectory directly, ResAD obtains an inertial reference (green arrow) as a counterfactual baseline. This forces the model to learn not what to do, but why it must deviate from this baseline, effectively linking actions (i.e., residuals) to their causal sources, like obstacles, rather than to statistical correlations.
@@ 1 Introduction
Conventional autonomous driving systems rely on a modular pipeline of perception, prediction, and planning components [26, 29, 7]. This cascaded design is prone to error propagation, leading to suboptimal or unsafe driving. In response to these limitations, End-to-End Autonomous Driving (E2EAD) has emerged as a compelling alternative [12, 2]. E2EAD reframes the driving problem by learning a direct mapping from raw sensor inputs to a future driving trajectory, from which control commands are derived, all within a single, unified framework [15, 33, 4, 10].
Recent years have witnessed extensive research into E2EAD methods, focusing on developing more effective representations [17, 3, 32], enhancing sensor fusion techniques [13, 37, 5, 28], and designing advanced architectures [35, 16, 31]. However, these existing methods are all attempting to answer the same question: “What is the future trajectory?” We argue this approach is inherently limited. As shown in Fig.1(a), the raw trajectory data exhibits a spatio-temporal non-uniformity, which leads to two critical issues that hinder real-world reliability and safety: Causal Confusion and the Planning Horizon Dilemma.
@@
Causal Confusion. The immense burden of mapping high-dimensional sensor data directly to a trajectory may cause the model to find shortcuts, relying on spurious correlations instead of the underlying causal logic governing safe driving [14, 20, 24]. For instance, a model might learn to associate braking with a lead vehicle’s brake lights but fail to understand the red light causing that vehicle to stop, leading it to dangerously follow the car through an intersection. Planning Horizon Dilemma. Trajectory data becomes more uncertain over longer horizons. Consequently, predictions for these distant waypoints often diverge significantly from the eventual ground truth, resulting in large loss values during training [2, 18]. This skews the optimization process, forcing the model to prioritize large, unpredictable long-term errors over the precision of the critical near-term path essential for immediate collision avoidance.
@@
To address these challenges, we propose  ResAD, a Normalized Residual Trajectory Modeling framework for End-to-End Autonomous Driving. As shown in Fig. 1 (b), our core idea is to decompose the complex prediction task into two distinct components: (1) a deterministic physics-based baseline, the inertial reference, obtained by extrapolating the vehicle’s current state to represent its default trajectory in the absence of active control; and (2) a learned residual, representing the necessary deviations from the inertial reference. By focusing specifically on deviations rather than the entire trajectory, ResAD shifts the learning objective from “What is the future trajectory?” to “Why must the trajectory change?”. This shift encourages the model to understand underlying causal factors (e.g., traffic rules, obstacles) instead of exploiting spurious correlations. To further mitigate the adverse effects of spatial scale variations during optimization, we conduct Point-wise Residual Normalization on the residuals. This technique prevents high-magnitude residuals at certain trajectory points from dominating the learning signal, ensuring that numerically small yet critically important adjustments are properly captured. Additionally, we strategically perturb the ego-vehicle’s state, generating diverse inertial references to counteract planning errors arising from sensor inaccuracies and guide the model toward a broader spectrum of high-quality trajectories. By embedding the fundamental physical prior of inertia into the model’s architecture, ResAD significantly simplifies the learning task, enabling more nuanced and precise driving behaviors. In summary, our contributions are as follows:
@@
  • We revisit the future-trajectory-prediction paradigm in E2EAD, and contend that the spatio-temporal non-uniformity of raw trajectory data leads to causal confusion and the planning horizon dilemma. This encourages a paradigm shift, moving from predicting the trajectory itself to modeling the reasons for its deviation.
@@
  • We propose ResAD, an E2EAD framework utilizing the Normalized Residual Trajectory Modeling. It first obtains an inertial reference by extrapolating the vehicle’s current state and then learns to predict the residual, i.e., the necessary deviations, relative to it. We further apply Point-wise Residual Normalization to the residuals, which prevents the optimization process from being dominated by long-horizon uncertainties.
@@
  • Extensive experiments and analyses validate the effectiveness of the proposed ResAD. On the NAVSIM benchmark, our method achieves state-of-the-art performance with scores of 88.6 for PDMS and 85.5 for EPDMS.
@@ 2.1 End-to-End Autonomous Driving
End-to-end autonomous driving (E2EAD) seeks to overcome the limitations of traditional modular pipelines, such as error accumulation and inter-module information loss [14, 31, 36]. Pioneering works like UniAD [12] introduced a planning-oriented architecture that jointly optimizes perception and forecasting to mitigate error propagation. VAD [17] further streamlined the pipeline with a fully vectorized scene representation, enabling the enforcement of explicit, instance-level safety constraints. More recently, generative models have become a new frontier in E2EAD research [6, 39, 27]. GoalFlow [34] introduces a goal-conditioned generative model that first selects an optimal goal point based on scene context and then uses Flow Matching to efficiently generate high-quality trajectories towards it. Despite these advances, existing methods predominantly rely on the direct prediction of future trajectories. In this work, we depart from this paradigm and introduce Normalized Residual Trajectory Modeling. Our method formulates a trajectory by decomposing it into a physically-based inertial reference and a learnable residual, offering a more structured and interpretable approach to trajectory representation.
Figure 2: The proposed ResAD framework. Multi-view images and LiDAR data are first processed and fused by a feature interaction encoder. We generate an inertial reference from the ego-vehicle’s state and perturb it into a cluster to ensure robustness and enable multi-modal predictions. Finally, diffusion decoders, conditioned on this reference cluster, merge the encoded features via cross-attention to output the planned trajectories.
@@ 2.2 Multimodal Planning
Most E2EAD systems produce a single, deterministic trajectory, an approach that struggles with the inherent diversity of real-world driving scenarios. To address this, several works have explored multimodal planning. VADv2 [3] proposes a probabilistic planning framework that outputs a distribution of future trajectories, which can be sampled to produce a diverse set of behaviors. The Hydra-MDP series [23, 21] employs policy distillation to select multiple trajectories from a vocabulary guided by an expert. GTRS [25] adopts a different strategy, scoring a set of pre-generated trajectories to ensure both diversity and safety. DiffusionDrive [27] highlights the challenge of mode collapse in generative-based models, addressing it by anchoring trajectory generation to a fixed cluster vocabulary. However, these methods fundamentally rely on a static, predefined vocabulary. This makes them both inefficient and restrictive, forcing them to evaluate irrelevant options while being unable to generate truly optimal trajectories outside the discrete set. Differently,  ResAD benefits from the unique trajectory modeling strategy that enables it to directly denoise from the Gaussian noise, yielding superior, context-aware multimodal trajectories.
@@ 3.1 Preliminaries
E2EAD aims to learn a unified policy, $\pi$, that directly maps raw sensor inputs, $\mathcal{O}$, to a sequence of waypoints, $\tau=\{(x_{t},y_{t})\}_{t=1}^{T_{f}}$, where $T_{f}$ denotes the planning horizon, and $(x_{t},y_{t})$ is the predicted future location of each waypoint at time $t$. We construct the proposed ResAD using a vanilla diffusion [11, 30] framework. The diffusion model defines a Markovian chain of diffusion forward process $q$ by gradually adding noise to sample data $z_{0}$, over a series of $T$ timesteps, which can be formulated as:
@@
where $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}.$ The hyperparameters $\beta_{t}$ controlling the amount of noise added at each step. As $t\xrightarrow{}T$, $z_{T}$ approaches a pure Gaussian noise distribution. We train the denoise model $\pi_{\theta}$ to predict $z_{0}$ from $z_{i}$ with the guidance of conditional information $c$, where $\theta$ are the trainable parameters. At the inference stage, the trained network is used to iteratively denoise a pure noise $z_{T}\sim\mathcal{N}(0,\mathbf{I})$ to produce a clean data sample $x_{0}$, which is defined as:
@@
In this work, we aim to solve the E2EAD via the diffusion model. Instead of directly generating the future trajectory points, we define the data samples as a set of normalized residuals.
@@ 3.2 Normalized Residual Trajectory Modeling
As illustrated in Fig. 2, ResAD takes multi-view images and LiDAR point clouds as input, which are fused by a Transfuser-style encoder. From the ego-vehicle state, we generate an inertial reference.  ResAD then perturbs this reference into a cluster to ensure robustness to state noise and to enable multi-modal predictions. The Diffusion Decoders employ cross-attention to merge the encoded features, using the inertial reference cluster as a condition to guide the training.
Trajectory Residual Modeling. The core idea of ResAD is to reframe trajectory prediction as a simpler, more interpretable learning problem. Instead of predicting the entire future trajectory from scratch, the model learns to predict the necessary correction to a simple, physics-based baseline. This baseline is an inertial reference trajectory, extrapolated from the ego-vehicle’s current status using a constant velocity model. This reframing compels the model to learn the control interventions required to deviate from the default path, focusing its capacity on the causal elements of driving.
@@
Point-wise Residual Normalization. A key challenge in trajectory prediction is the scale variance of coordinates across the time horizon. Points further in the future have numerically larger values, which can cause the optimization to be dominated by far-field errors, neglecting the fine-grained, safety-critical adjustments required in the near field. As shown in Fig. 1(a), while residual modeling mitigates this by focusing on deviations, the scale issue within the residuals themselves persists. We propose Point-wise Residual Normalization (PRNorm) to resolve this.
@@
“C”: Camera, “L”: LiDAR. The best and second-best scores are highlighted in bold and underlined, respectively. * For fair comparison, we use the official scores of versions with the same backbone.
Inertia Reference Perturbation. Driving is an inherently multi-modal task. Most methods depend on a fixed trajectory vocabulary, where most options are irrelevant to the current scene, causing inefficiency. ResAD, circumvents this issue through Trajectory Residual Modeling, which generates multi-modal trajectories by perturbing its Inertial Reference. This approach is doubly beneficial. It forces the model to learn resilience against noise from ego-sensors like GPS and IMU. On the other hand, it creates a set of intent hypotheses by generating a cluster of slightly varied inertial references. The network then produces a full trajectory for each hypothesis, naturally yielding a diverse set of context-relevant paths.
@@
During inference, the denoising process starts with $K_{\mathrm{infer}}$ Gaussian noise to generate residuals. We take 2 timesteps to get the final predictions $\{\hat{\bm{r}}_{k}\}_{k=1}^{K_{\mathrm{infer}}}$ by DDIM [30]. Then the predicted residuals are added to the corresponding perturbed inertia reference to get the multi-modal trajectory $\{\hat{\tau}_{k}\}_{k=1}^{K_{\mathrm{infer}}}$.

### v2

@@ Abstract
End-to-end autonomous driving (E2EAD) systems, which learn to predict future trajectories directly from sensor data, are fundamentally challenged by the inherent spatio-temporal imbalance of trajectory data. This imbalance creates a significant optimization burden, causing models to learn spurious correlations instead of robust driving logic, while also prioritizing uncertain, distant predictions, thereby compromising immediate safety. To address these issues, we propose ResAD, a novel Normalized Residual Trajectory Modeling framework. Instead of predicting the future trajectory directly, our approach reframes and simplifies the learning task by predicting the residual deviation from a deterministic inertial reference. This inertial reference serves as a strong physical prior, compelling the model to move beyond simple pattern-matching and instead focus its capacity on learning the necessary, context-driven deviations (e.g., traffic rules, obstacles) from this default, inertially-guided path. To mitigate the optimization imbalance caused by uncertain, long-term horizons, ResAD further incorporates Point-wise Normalization of the predicted residual. This technique re-weights the optimization objective, preventing large-magnitude errors associated with distant, uncertain waypoints from dominating the learning signal. On the NAVSIM v1 and v2 benchmarks, ResAD achieves state-of-the-art results of 88.8 PDMS and 85.5 EPDMS with only two denoising steps, demonstrating that ResAD significantly simplifies the learning task and improves planning performance. The code will be released to facilitate further research.
@@
Figure 1: The core motivation for ResAD: Addressing the Challenge of Imbalanced Trajectory Data. (a) Visualization of the longitudinal data distribution of dataset trajectories under three modeling strategies. Raw Trajectories exhibit significant mean drift and increasing variance, creating a planning horizon dilemma. Our Trajectory Residual centers the distribution, and the Normalized Residual further stabilizes the variance for a simpler, balanced learning objective. (b) Conceptual comparison. Existing methods learn complex raw trajectories directly, risking reliance on spurious correlations. Our ResAD simplifies the task by learning to predict only the necessary residual deviation from a strong physical prior (the inertial reference), focusing the model on context-aware corrections.
@@ 1 Introduction
Conventional autonomous driving systems rely on a modular pipeline of perception, prediction, and planning components [liang2020pnpnet, luo2018fast, chitta2022transfuser, wang2022ltp, wu2022trajectory]. This cascaded design is prone to error propagation, leading to suboptimal or unsafe driving. In response to these limitations, End-to-End Autonomous Driving (E2EAD) has emerged as a compelling alternative [hu2023uniad, chen2024end, li2024ego, sima2024drivelm]. E2EAD reframes the driving problem by learning a direct mapping from raw sensor inputs to a future driving trajectory, from which control commands are derived, all within a single, unified framework [jia2023thinktwice, weng2024para, zhou2025opendrivevla, hamdan2025eta].
@@
Recent years have witnessed extensive research into E2EAD methods, focusing on developing more effective representations [jiang2023vad, chen2024vadv2, sun2025sparsedrive, zhang2025bridging], enhancing sensor fusion techniques [Jaeger2023ICCV, ye2023fusionad, chen2024dualat, liu2025gaussianfusion], and designing advanced architectures [yang2024visual, jia2025drivetransformer, song2025don, RAD, zhengdiffusion]. However, these existing methods are all attempting to answer the same question: “What is the future trajectory?” We argue this approach is inherently limited. As shown in Fig.1 (a), the raw trajectory data exhibits a spatio-temporal non-uniformity, which leads to two critical issues that hinder real-world reliability and safety: (1) Spurious Correlations. The immense burden of mapping high-dimensional sensor data directly to a trajectory may cause the model to find shortcuts, relying on spurious correlations instead of the underlying robust logic governing safe driving [jia2023driveadapter, li2024exploring, shao2023reasonnet]. For instance, a model might learn to associate braking with a lead vehicle’s brake lights but fail to understand the red light causing that vehicle to stop, leading it to dangerously follow the car through an intersection. (2) Planning Horizon Dilemma. Trajectory data becomes more uncertain over longer horizons. Consequently, predictions for these distant waypoints often diverge significantly from the eventual ground truth, leading to large loss values during training [salzmann2020trajectron++, veer2022receding, chen2024end, madjid2025trajectory, jiang2025survey]. This skews the optimization process, forcing the model to prioritize large, unpredictable long-term errors over the precision of the critical near-term path essential for immediate collision avoidance.
@@
To address these challenges, we propose  ResAD, a Normalized Residual Trajectory Modeling framework for End-to-End Autonomous Driving. As shown in Fig. 1 (b), our core idea is to decompose the complex prediction task into two distinct components: (1) a deterministic physics-based baseline, the inertial reference, representing the vehicle’s default path in the absence of active control; and (2) a learned residual, the necessary deviations from the inertial reference. By focusing specifically on deviations rather than the entire trajectory, ResAD reframes the learning objective from “What is the future trajectory?” to “Why must the trajectory change?”.This shift significantly simplifies the learning problem. By providing a strong physical prior, the model can focus its capacity on learning the mapping from sensor inputs (e.g., traffic rules, obstacles) to the required corrections (i.e., residuals), rather than expending it on learning complex spatio-temporal dynamics from scratch. To further mitigate the adverse effects of spatial scale variations during optimization, we introduce Point-wise Residual Normalization for the residuals. This technique prevents high-magnitude residuals at certain trajectory points from dominating the learning signal, ensuring that numerically small yet critically important adjustments are properly captured. Additionally, we strategically perturb the ego-vehicle’s state, generating diverse inertial references to counteract planning errors arising from sensor inaccuracies and guide the model toward a broader spectrum of high-quality trajectories. In summary, our contributions are as follows:
@@
  • We revisit the future-trajectory-prediction paradigm in E2EAD, and contend that the spatio-temporal non-uniformity of raw trajectory data leads to spurious correlations and the planning horizon dilemma.
@@
  • We propose ResAD, an E2EAD framework based on the Normalized Residual Trajectory Modeling. It simplifies the learning task by first obtaining an inertial reference and then learning to predict the necessary residual deviations relative to it. We further introduce Point-wise Residual Normalization to mitigate optimization imbalance caused by long-horizon uncertainties.
@@
  • Extensive experiments and analyses validate the effectiveness of the proposed ResAD. On the NAVSIM v1 and v2 benchmark, our method achieves state-of-the-art performance with scores of 88.8 for PDMS and 85.5 for EPDMS.
@@ 2.1 End-to-End Autonomous Driving
End-to-end autonomous driving (E2EAD) seeks to overcome the limitations of traditional modular pipelines, such as error accumulation and inter-module information loss [jia2023driveadapter, min2024driveworld, hanselmann2022king, nie2024reason2drive, jia2024bench2drive, song2025don, sun2025generalizing]. Pioneering works like UniAD [hu2023uniad] introduced a planning-oriented architecture that jointly optimizes perception and forecasting to mitigate error propagation. VAD [jiang2023vad] further streamlined the pipeline with a fully vectorized scene representation, enabling the enforcement of explicit, instance-level safety constraints. More recently, generative models have become a new frontier in E2EAD research [chi2023diffusion, chen2024ppad, feng2025artemis, zheng2024genad, liao2025diffusiondrive, xu2025knowledge, jiang2025diffvla, li2025discrete]. GoalFlow [xing2025goalflow] introduces a goal-conditioned generative model that first selects an optimal goal point based on scene context and then uses Flow Matching to efficiently generate high-quality trajectories towards it. Despite these advances, existing methods predominantly rely on the direct prediction of future trajectories. In this work, we depart from this paradigm by introducing Normalized Residual Trajectory Modeling. Our method formulates a trajectory by decomposing it into a physics-based inertial reference and a learnable residual, offering a more structured and interpretable approach to trajectory representation.
Figure 2: The proposed ResAD framework. Instead of predicting the entire trajectory, ResAD establishes a strong physical prior based on the vehicle’s current state: the Inertial Reference. By applying Inertial Reference Perturbation, the framework generates a diverse set of initial intent hypotheses. Finally, the diffusion decoder (DiffDecoder), conditioned on these references, learns to predict the necessary Normalized Residuals. We highlight the 1st-ranked and 5th-ranked output trajectories, denote as Top-1 Traj and Top-5 Traj.
@@ 2.2 Multimodal Planning
Most E2EAD systems produce a single, deterministic trajectory, an approach that struggles with the inherent diversity of real-world driving scenarios. To address this, several works have explored multimodal planning [chen2024vadv2, li2025generalized, song2025breaking, guo2025ipad]. VADv2 [chen2024vadv2] proposes a probabilistic planning framework that outputs a distribution of future trajectories, which can be sampled to produce a diverse set of behaviors. The Hydra-MDP series [li2024hydra, li2025hydra++] employs policy distillation to select multiple trajectories from a vocabulary guided by an expert. GTRS [li2025generalizedtrajectoryscoringendtoend] adopts a different strategy, scoring a set of pre-generated trajectories to ensure both diversity and safety. DiffusionDrive [liao2025diffusiondrive] highlights the challenge of mode collapse in generative-based models, addressing it by anchoring trajectory generation to a fixed cluster vocabulary. However, these methods fundamentally rely on a static, predefined vocabulary. This makes them both inefficient and restrictive, forcing them to evaluate irrelevant options while being unable to generate truly optimal trajectories outside the discrete set. In contrast,  ResAD benefits from its unique trajectory modeling strategy, which enables it to directly denoise from the Gaussian noise. This approach yields superior, context-aware multimodal trajectories without relying on a fixed vocabulary.
@@ 3.1 Preliminaries
E2EAD aims to learn a unified policy, $\pi$, that directly maps raw sensor inputs, $\mathcal{O}$, to a sequence of waypoints, $\tau=\{(x_{t},y_{t})\}_{t=1}^{T_{f}}$, where $T_{f}$ denotes the planning horizon, and $(x_{t},y_{t})$ is the predicted future location of each waypoint at time $t$. We construct the proposed ResAD using a vanilla diffusion [ho2020denoising, songdenoising] framework. The diffusion model defines a Markovian chain of diffusion forward process $q$ by gradually adding noise to sample data $z_{0}$, over a series of $T$ timesteps, which can be formulated as:
@@
where $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}.$ The hyperparameters $\beta_{t}$ controlling the amount of noise added at each step. As $t\xrightarrow{}T$, $z_{T}$ approaches a pure Gaussian noise distribution. We train the denoise model $\pi_{\theta}$ to predict $z_{0}$ from $z_{i}$ with the guidance of conditional information $c$, where $\theta$ are the trainable parameters. At the inference stage, the trained network is used to iteratively denoise a pure noise $z_{T}\sim\mathcal{N}(0,\mathbf{I})$ to produce a clean data sample $z_{0}$, which is defined as:
@@
In this work, instead of directly generating the future trajectory points, we define the data samples as a set of normalized residuals.
@@ 3.2 Normalized Residual Trajectory Modeling
As illustrated in Fig. 2, ResAD takes multi-view images and LiDAR point clouds as input, which are fused by a Transfuser-style encoder [chitta2022transfuser]. We first generate an inertial reference from the ego-vehicle’s current state.  ResAD then perturbs this reference to generate a cluster, ensuring robustness and enabling multimodal predictions. Finally, the Diffusion Decoders employ cross-attention to merge the encoded features, using the inertial reference cluster as a condition to guide the denoising process.
Trajectory Residual Modeling. The core idea of ResAD is to reframe the complex trajectory prediction task into a simpler, more interpretable learning problem. Instead of predicting the entire future trajectory from scratch, we use a constant velocity mode [rhinehart2018r2p2] to extrapolate the inertial reference trajectory from the ego-vehicle’s current state, and predict the residual. This remodeling compels the model to learn the control interventions required to deviate from the default reference, focusing on the critical, high-level driving decisions needed by the scene.
@@
Point-wise Residual Normalization. A key challenge in planning is the scale variance of coordinates across the time horizon. Points further in the future have numerically larger values, which can cause the optimization to be dominated by far-field errors, neglecting the fine-grained, safety-critical adjustments required in the near field. As shown in Fig. 1(a), while residual modeling mitigates this by focusing on deviations, the scale issue within the residuals themselves persists. We propose Point-wise Residual Normalization (PRNorm) to resolve this.
@@
[…]
“C”: Camera, “L”: LiDAR. The best and second-best scores are highlighted in bold and underlined, respectively.
Inertial Reference Perturbation. Driving is an inherently multi-modal task. Most methods depend on a fixed trajectory vocabulary, where most options are irrelevant to the current scene, causing inefficiency. ResAD generates multimodal trajectories by perturbing Inertial Reference. This approach offers a dual benefit. First, it generates diverse intent hypotheses by creating varied inertial references and corresponding residuals. Second, it compels the model to handle noise inherent in ego-sensors such as GPS and IMU.
@@
During inference, the denoising process starts with $K_{\mathrm{infer}}$ Gaussian noise to generate residuals. We take 2 timesteps to get the final predictions $\{\hat{\bm{r}}_{k}\}_{k=1}^{K_{\mathrm{infer}}}$ by DDIM [songdenoising]. Then the predicted residuals are added to the corresponding perturbed inertial reference to get the multi-modal trajectory $\{\hat{\tau}_{k}\}_{k=1}^{K_{\mathrm{infer}}}$.